OpenStudy (anonymous):

How do I factor 16x^2 -49

OpenStudy (haseeb96):

very easy 16x^2-49=0 16x^2 =49 x^2 = 49/16 now taking square root on both sides u will get x = 7/4 and x =-7/4

OpenStudy (anonymous):

It needs to look like (x +/- _ )(x+/- _ )